# Polarized and unpolarized gluon PDFs: generative machine learning applications for lattice QCD matrix elements at short distance and large momentum

### v1

@@ V.1 ITD for polarized gluon distribution
In Fig. 12, we present the $\Delta\mathfrak{M}_{g}$ correlation function for the gluon helicity distribution. The contamination-free $\Delta{\mathfrak{M}_{g}}(\omega)$ can now be matched to the light cone $\Delta{\mathcal{I}}_{g}(\omega,\mu)$ and the singlet quark ITD $\Delta{\mathcal{I}}_{S}(\omega,\mu)$ in the $\overline{\rm{MS}}$ scheme using the factorization relation [39] up to power corrections,
@@
where $N_{c}=3$, $\bar{u}\equiv(1-u)$, $\gamma_{E}$ is the Euler–Mascheroni constant, $\Delta{\mathcal{B}}_{gq}=1-(1-u)^{2}$, and $\langle x\rangle_{g}(\mu=2~{}{\rm GeV})$ is chosen to be $0.427$ from [66]. We choose $z=2a$ and $\mu=2$ GeV in the matching Eq. (V.1) and ignore the effect of the singlet-quark contributions (which requires a separate LQCD calculation and is a subject for future investigations). Additionally, varying values of $z$ or $\mu$ have minimal effects on the matched $\Delta{\mathcal{I}}_{g}$ within the current statistical uncertainty due to the large uncertainty of the present LQCD matrix elements [37].
In Fig. 13, we present the $\Delta{\mathcal{I}}_{g}(\omega)$ in the $\overline{\rm{MS}}$ scheme determined from the above-mentioned ML algorithms. We compare the results with the phenomenological ITD constructed from the NNPDF3.1 global fit [53]. As described above, the RF model produces the best result.
Figure 13: (Upper panel) Polarized gluon light-cone Ioffe-time distributions from various generative models of machine learning after removing contamination term $\frac{m_{p}^{2}}{p_{z}^{2}}\omega\Delta{\mathcal{M}}_{pp}(\omega)$ from the LQCD matrix elements $\Delta\mathfrak{M}(\omega)$ using RF, GBT, and XGB machine learning algorithm and the implementation of the perturbative matching kernel in the $\overline{\rm{MS}}$ renormalization scheme at 2 GeV. $\Delta\mathfrak{M}_{g}(\omega)$ has been normalized by the gluon momentum fraction $\langle x\rangle_{g}(\mu)$ before converting to $\Delta\mathcal{I}_{g}(\omega,\mu)$. (Lower panel) We compare our results to phenomenological ITD constructed from the polarized gluon PDFs in NNPDF3.1 global fit [53].
@@ V.2 ITD for the unpolarized gluon distribution
In order to determine the ITD in the $\overline{\rm{MS}}$-scheme, we use the one-loop matching relation [38],
@@
where we have neglected the quark-gluon mixing. $\mathcal{I}_{g}(0,\mu^{2})$ is the gluon momentum fraction $\langle x\rangle_{g}(\mu)$ and $B_{gg}(u)$ is the Altarelli-Parisi kernel. In the left panel of Fig. 15, we present the unpolarized gluon ITDs derived from ML-generated lattice data at $p_{z}=2.05$ GeV and compare these with the phenomenological ITD extracted from the NNPDF3.1 [53] in the $\overline{\rm{MS}}$ scheme at $\mu=2$ GeV. A similar comparison is made with the data lattice at $p_{z}=2.46$ GeV. In contrast to the polarized ITD, where the momentum is very large and thus the $z$-values are very small for $\omega=14$, for the unpolarized ITD with $p_{z}=2.46$ GeV, $\omega=14$ corresponds to $z\approx 1$ fm. In future, with more precise gluonic matrix elements, we aim to isolate the $z$-dependence in the LQCD data and apply ML techniques using both $z$ and $p_{z}$ as input features. This will allow us to determine the unpolarized gluon ITD at short distances and higher momenta. Unfortunately, current statistics do not permit us to perform such ML analyses.
@@
  • The ML-generated ITD up to $\omega\sim 14$ will be significantly more important for reconstructing PDFs compared to the ITD limited to $\omega\approx 7$ in [31].
@@
Figure 15: Unpolarized gluon Ioffe-time distributions from various generative models of machine learning after implementing the perturbative matching kernel in the $\overline{\rm{MS}}$ renormalization scheme at $\mu$ = 2 GeV. We compare our results with the phenomenological ITD constructed from the unpolarized gluon PDF in NNPDF3.1 global fit [53].
@@ VI Neural network reconstruction of PDFs
To extract the polarized gluon PDF $x\Delta g(x,\mu)$ from the ITDs presented in the previous section, we need to solve the following inverse problem [18]:
@@
We determine the $x\Delta g(x)$ and $xg(x)$ distributions in Eqs. (17) and (18) using the similar neural network (NN) architecture used in [37]. Instead of directly parametrizing the distribution function by a NN, we take a prefit function for the polarized gluon PDF from one of the parametrizations in [43],
@@
The fitting procedure is to minimize the loss function defined as the $\chi^{2}$ between the NN predictions and the data points. In each fit, we randomly select 80% of the data points generated by the ML algorithms in the previous sections for $\omega\leq 10,12,14$ and leave the remaining 20% of the data points in the validation sample to evaluate the performance of the model on unseen data. To keep the possibility of finding multiple minima, the initial parameters of the NN are randomly generated. The loss value of the full dataset is monitored during the training. It generally decreases at the beginning and starts to increase when overfitting happens, with small fluctuations from epoch to epoch all the time. To prevent the accidental selection of small loss value points, some early epochs are eliminated. We stop the training process when there is no improvement in the total loss value for $3000$ epochs and revert to the best result obtained. The result from the epoch with the least total loss function is saved. The results always converge to the same region. In addition, the partition of the training sample and the validation sample was not fixed from time to time, and the result does not show dependence on the partition within the uncertainties inheriting from the LQCD data. As has been investigated in [37], the systematic uncertainty from the choice of the NN is negligible and the uncertainty is dominated by the lattice data. One may perform the analysis using an NN with an additional hidden layer, i.e. an NN with four hidden layers containing 32, 32, 32, and 8 neurons respectively as more hidden layers mean more flexible parametrization of the function. We find almost no noticeable effects on the resulting distribution function with the uncertainty. One can also modify the activation functions of the layers, e.g. the exponential linear unit function and the hyperbolic tangent function. Once again, similar outcomes of the PDFs are obtained. Since these modifications of the NN have already been investigated and numerically demonstrated in [37], we do not repeat the calculations here. Rather, we concentrate on the PDF reconstructions for different ranges of $\omega\in[10,14]$ which have the largest effects on the resulting PDFs.
Based on the outcomes in Sec. IV, we consider the ITDs from the two best fits, from RF and GBT. In Fig. 16, we present $x\Delta g(x)$ reconstructed from the ITD data up to $\omega_{\rm max}=10,12,14$ and compare them with that presented in [53].
Figure 16: Neural network reconstruction of the polarized gluon PDF $x\Delta g(x)$ from the ITDs generated with RF and GBT machine learning algorithms. The neural network reconstructions of the PDFs have been done for three different values of $\omega_{\rm max}=10,12,14$ of the ITD, $\Delta\mathcal{I}_{g}(\omega,\mu)$. The grey band is obtained from the NNPDF global analysis in [53].
@@
Similarly, for the unpolarized gluon distribution, we take the prefit function from a fit in [43],
@@
In contrast, as shown in the PDF reconstruction in Fig. 18, it is immediately evident that the ITD with $\omega_{\rm max}=14$ is much more effective for PDF reconstruction compared to the ITD with $\omega_{\rm max}=7$ in constraining uncertainties in the mid- and large-$x$ regions of the constructed PDF. Specifically, the LQCD data is expected to be most sensitive to the PDF’s mid-$x$ region, where significantly smaller uncertainties can be achieved with the ITD with $\omega_{\rm max}=14$ compared to the ITD with $\omega_{\rm max}=7$. Another significant achievement is that the ML-generated ITD can reduce uncertainties in the mid to large $x\sim 0.9$ region. A precise estimate of $xg(x)$ at large $x$ is crucial for generating accurate predictions of both signal and background in searches for new massive particles at the LHC [68]. Since there is a difference between the ITD from the NNPDF dataset and ITD from the LQCD calculation at an unphysical pion mass and a coarse lattice spacing (see Fig. 15), we expect the reconstructed PDF to differ from the NNPDF. The major advantage highlighted here is that with the LQCD and the ML-generated data up to $\omega=14$, the impact on constraining the $xg(x)$ distribution at mid to large $x$ can be significant.
Figure 17: Neural network reconstruction of the unpolarized gluon PDF $xg(x)$ from the ITDs generated with RF and GBT machine learning algorithms. The neural network reconstructions of the PDFs have been done for three different values of $\omega_{\rm max}=10,12,14$ of the ITD, $\mathcal{I}_{g}(\omega,\mu)$. The grey band is obtained from the NNPDF global analysis in [53].
Figure 18: Neural network reconstruction of the unpolarized gluon PDF using the RF-generated data for $\omega_{\rm max}=7$ and $14$. We compare our results with the gluon PDFs extracted from NNPDF3.1 global fit [53]. The prefit function (22) is also plotted.
@@ VII Impacts of the polarized and unpolarized gluon distribution results
In the first LQCD determination of the gluon helicity PDF [37], it was predicted that the gluon helicity PDF is positive in the mid- to large-$x$ region within uncertainty. This contrasts with two global analyses [69, 70], which reported that both positive and negative solutions for $x\Delta g(x)$ were equally capable of describing the experimental data. Subsequently, a similar conclusion from [37] was also found in [71] with the fundamental requirement that the physical cross-sections must not be negative. Following [71], the analyses in [69, 70] were revisited in [72], and a similar observation of the positivity of $x\Delta g(x)$ was found. In the present work, using the physics-informed generative ML algorithms, we have solidified our previous results in [37] with higher precision and up to a larger value of $\omega\sim 14$. While caution is needed regarding the sign of $x\Delta g(x)$ in the small-$x$ region, where LQCD data is not sensitive to $x\Delta g(x)$, the PDFs determined in Fig. 16 show a positive distribution in the interval $x\in[0.05,0.95]$. A recent phenomenological calculation [73] has also determined $x\Delta g(x)$ to be positive. In fact, this finding is also consistent with the LQCD calculation [74] at the physical pion mass, continuum, and infinite volume limits. The LQCD calculation in [74] obtained the $x$-integrated value of $\Delta G=0.251(47)(16)$ using a local matrix element [59].
@@
As noted in the previous section, the $xg(x)$ falls off much more slowly than the NNPDF result at large $x$. This may be due to the LQCD calculation being performed at a heavier pion mass and a coarse lattice spacing. Additionally, we observe that because of the large uncertainty in the LQCD matrix elements, the ML algorithms could not isolate any $z$-dependence from the LQCD matrix elements and used data up to $z=0.56$ fm (unlike the polarized data, where Eq. (7) required the use of the LQCD data up to $z=0.36$ fm). As discussed in Sec. V, the large $z$ data used in the ML-generated ITD might be affected by the higher-twist contaminations, leading to a slower fall-off of the PDF. However, as mentioned earlier, the goal of this work is to demonstrate how ML can open an avenue to constrain $xg(x)$ distribution in the mid-to-large $x$ region using LQCD calculations. Future precise LQCD matrix elements will enable isolating the $z$-dependence as an input feature in the ML analyses to obtain $xg(x)$ in the physical and continuum limits. In addition, physics-informed ML application on the LQCD data can have a significant impact on extracting the gluon generalized parton distributions (GPDs). These prospects have been discussed in [75, 76, 77].
@@ VIII Conlusion and outlook
This work also demonstrates that LQCD nonlocal spatial matrix elements for the gluon helciity distribution beyond $0.36$ fm may contain significant higher-twist contamination, which can impact the PDF determination and should be avoided to ensure the applicability of factorizing the LQCD matrix elements into perturbative Wilson coefficients and nonperturbative PDFs with controllable power corrections. To address this issue, one can train the ML algorithms to learn from short-distance correlations at $z\lesssim 0.36$ fm, thereby minimizing possible contamination from higher-twist effects $\mathcal{O}(\Lambda_{\rm QCD}^{2}z^{2})$ at large spatial distances for a successful reconstruction of the PDFs. Looking forward, the demonstrated interface between LQCD and ML holds promising potential for investigating the higher-twist contributions at the level of LQCD correlation functions when precise data are available using $z$, $p_{z}$, and $\omega$ as input features in the ML, an area that remains largely unexplored.

### v2

@@ V.1 ITD for polarized gluon distribution
In Fig. 14, we present the $\Delta\mathfrak{M}_{g}$ correlation function for the gluon helicity distribution. The contamination-free $\Delta{\mathfrak{M}_{g}}(\omega)$ can now be matched to the light cone $\Delta{\mathcal{I}}_{g}(\omega,\mu)$ and the singlet quark ITD $\Delta{\mathcal{I}}_{S}(\omega,\mu)$ in the $\overline{\rm{MS}}$ scheme using the factorization relation Balitsky et al. (2022) up to power corrections,
@@
where $N_{c}=3$, $\bar{u}\equiv(1-u)$, $\gamma_{E}$ is the Euler–Mascheroni constant, $\Delta{\mathcal{B}}_{gq}=1-(1-u)^{2}$, and $\langle x\rangle_{g}(\mu=2~{}{\rm GeV})$ is chosen to be $0.427$ from Alexandrou et al. (2020). We choose $z=2a$ and $\mu=2$ GeV in the matching Eq. (V.1) and ignore the effect of the singlet-quark contributions (which requires a separate LQCD calculation and is a subject for future investigations). Additionally, varying values of $z$ or $\mu$ have minimal effects on the matched $\Delta{\mathcal{I}}_{g}$ within the current statistical uncertainty due to the large uncertainty of the present LQCD matrix elements Khan et al. (2023).
In Fig. 15, we present the $\Delta{\mathcal{I}}_{g}(\omega)$ in the $\overline{\rm{MS}}$ scheme determined from the above-mentioned ML algorithms. We compare the results with the phenomenological ITD constructed from the NNPDF3.1 global fit Ball, Richard D. et al. (2017). As described above, the RF model produces the best result.
Figure 15: (Upper panel) Polarized gluon light-cone Ioffe-time distributions from various generative models of machine learning after removing contamination term $\frac{m_{p}^{2}}{p_{z}^{2}}\omega\Delta{\mathcal{M}}_{pp}(\omega)$ from the LQCD matrix elements $\Delta\mathfrak{M}(\omega)$ using RF, GBT, and XGB machine learning algorithm and the implementation of the perturbative matching kernel in the $\overline{\rm{MS}}$ renormalization scheme at 2 GeV. $\Delta\mathfrak{M}_{g}(\omega)$ has been normalized by the gluon momentum fraction $\langle x\rangle_{g}(\mu)$ before converting to $\Delta\mathcal{I}_{g}(\omega,\mu)$. (Lower panel) We compare our results to phenomenological ITD constructed from the polarized gluon PDFs in NNPDF3.1 global fit Ball, Richard D. et al. (2017).
@@ V.2 ITD for the unpolarized gluon distribution
In order to determine the ITD in the $\overline{\rm{MS}}$-scheme, we use the one-loop matching relation Balitsky et al. (2020),
@@
where we have neglected the quark-gluon mixing. $\mathcal{I}_{g}(0,\mu^{2})$ is the gluon momentum fraction $\langle x\rangle_{g}(\mu)$ and $B_{gg}(u)$ is the Altarelli-Parisi kernel. In the left panel of Fig. 17, we present the unpolarized gluon ITDs derived from ML-generated lattice data at $p_{z}=2.05$ GeV and compare these with the phenomenological ITD extracted from the NNPDF3.1 Ball, Richard D. et al. (2017) in the $\overline{\rm{MS}}$ scheme at $\mu=2$ GeV. A similar comparison is made with the data lattice at $p_{z}=2.46$ GeV. In contrast to the polarized ITD, where the momentum is very large and thus the $z$-values are very small for $\omega=14$, for the unpolarized ITD with $p_{z}=2.46$ GeV, $\omega=14$ corresponds to $z\approx 1$ fm. In future, with more precise gluonic matrix elements, we aim to isolate the $z$-dependence in the LQCD data and apply ML techniques using both $z$ and $p_{z}$ as input features. This will allow us to determine the unpolarized gluon ITD at short distances and higher momenta. Unfortunately, current statistics do not permit us to perform such ML analyses.
@@
  • The ML-generated ITD up to $\omega\sim 14$ will be significantly more important for reconstructing PDFs compared to the ITD limited to $\omega\approx 7$ in Khan et al. (2021).
@@
Figure 17: Unpolarized gluon Ioffe-time distributions from various generative models of machine learning after implementing the perturbative matching kernel in the $\overline{\rm{MS}}$ renormalization scheme at $\mu$ = 2 GeV. We compare our results with the phenomenological ITD constructed from the unpolarized gluon PDF in NNPDF3.1 global fit Ball, Richard D. et al. (2017).
@@ VI Neural network reconstruction of PDFs
To extract the polarized gluon PDF $x\Delta g(x,\mu)$ from the ITDs presented in the previous section, we need to solve the following inverse problem Braun et al. (1995):
@@
We determine the $x\Delta g(x)$ and $xg(x)$ distributions in Eqs. (17) and (18) using the similar neural network (NN) architecture used in Khan et al. (2023). Instead of directly parametrizing the distribution function by a NN, we take a prefit function for the polarized gluon PDF from one of the parametrizations in Sufian et al. (2021),
@@
The fitting procedure is to minimize the loss function defined as the $\chi^{2}$ between the NN predictions and the data points. In each fit, we randomly select 80% of the data points generated by the ML algorithms in the previous sections for $\omega\leq 10,12,14$ and leave the remaining 20% of the data points in the validation sample to evaluate the performance of the model on unseen data. To keep the possibility of finding multiple minima, the initial parameters of the NN are randomly generated. The loss value of the full dataset is monitored during the training. It generally decreases at the beginning and starts to increase when overfitting happens, with small fluctuations from epoch to epoch all the time. To prevent the accidental selection of small loss value points, some early epochs are eliminated. We stop the training process when there is no improvement in the total loss value for $3000$ epochs and revert to the best result obtained. The result from the epoch with the least total loss function is saved. The results always converge to the same region. In addition, the partition of the training sample and the validation sample was not fixed from time to time, and the result does not show dependence on the partition within the uncertainties inheriting from the LQCD data. As has been investigated in Khan et al. (2023), the systematic uncertainty from the choice of the NN is negligible and the uncertainty is dominated by the lattice data. One may perform the analysis using an NN with an additional hidden layer, i.e. an NN with four hidden layers containing 32, 32, 32, and 8 neurons respectively as more hidden layers mean more flexible parametrization of the function. We find almost no noticeable effects on the resulting distribution function with the uncertainty. One can also modify the activation functions of the layers, e.g. the exponential linear unit function and the hyperbolic tangent function. Once again, similar outcomes of the PDFs are obtained. Since these modifications of the NN have already been investigated and numerically demonstrated in Khan et al. (2023), we do not repeat the calculations here. Rather, we concentrate on the PDF reconstructions for different ranges of $\omega\in[10,14]$ which have the largest effects on the resulting PDFs.
Based on the outcomes in Sec. IV, we consider the ITDs from the two best fits, from RF and GBT. In Fig. 18, we present $x\Delta g(x)$ reconstructed from the ITD data up to $\omega_{\rm max}=10,12,14$ and compare them with that presented in Ball, Richard D. et al. (2017).
Figure 18: Neural network reconstruction of the polarized gluon PDF $x\Delta g(x)$ from the ITDs generated with RF and GBT machine learning algorithms. The neural network reconstructions of the PDFs have been done for three different values of $\omega_{\rm max}=10,12,14$ of the ITD, $\Delta\mathcal{I}_{g}(\omega,\mu)$. The grey band is obtained from the NNPDF global analysis in Ball, Richard D. et al. (2017).
@@
Similarly, for the unpolarized gluon distribution, we take the prefit function from a fit in Sufian et al. (2021),
@@
In contrast, as shown in the PDF reconstruction in Fig. 20, it is immediately evident that the ITD with $\omega_{\rm max}=14$ is much more effective for PDF reconstruction compared to the ITD with $\omega_{\rm max}=7$ in constraining uncertainties in the mid- and large-$x$ regions of the constructed PDF. Specifically, the LQCD data is expected to be most sensitive to the PDF’s mid-$x$ region, where significantly smaller uncertainties can be achieved with the ITD with $\omega_{\rm max}=14$ compared to the ITD with $\omega_{\rm max}=7$. Another significant achievement is that the ML-generated ITD can reduce uncertainties in the mid to large $x\sim 0.9$ region. A precise estimate of $xg(x)$ at large $x$ is crucial for generating accurate predictions of both signal and background in searches for new massive particles at the LHC Nocera and Ubiali (2018). Since there is a difference between the ITD from the NNPDF dataset and ITD from the LQCD calculation at an unphysical pion mass and a coarse lattice spacing (see Fig. 17), we expect the reconstructed PDF to differ from the NNPDF. The major advantage highlighted here is that with the LQCD and the ML-generated data up to $\omega=14$, the impact on constraining the $xg(x)$ distribution at mid to large $x$ can be significant.
Figure 19: Neural network reconstruction of the unpolarized gluon PDF $xg(x)$ from the ITDs generated with RF and GBT machine learning algorithms. The neural network reconstructions of the PDFs have been done for three different values of $\omega_{\rm max}=10,12,14$ of the ITD, $\mathcal{I}_{g}(\omega,\mu)$. The grey band is obtained from the NNPDF global analysis in Ball, Richard D. et al. (2017).
Figure 20: Neural network reconstruction of the unpolarized gluon PDF using the RF-generated data for $\omega_{\rm max}=7$ and $14$. We compare our results with the gluon PDFs extracted from NNPDF3.1 global fit Ball, Richard D. et al. (2017). The prefit function (22) is also plotted.
@@ VII Impacts of the polarized and unpolarized gluon distribution results
In the first LQCD determination of the gluon helicity PDF Khan et al. (2023), it was predicted that the gluon helicity PDF is positive in the mid- to large-$x$ region within uncertainty. This contrasts with two global analyses Zhou et al. (2022); Whitehill et al. (2022), which reported that both positive and negative solutions for $x\Delta g(x)$ were equally capable of describing the experimental data. Subsequently, a similar conclusion from Khan et al. (2023) was also found in de Florian et al. (2024) with the fundamental requirement that the physical cross-sections must not be negative. Following de Florian et al. (2024), the analyses in Zhou et al. (2022); Whitehill et al. (2022) were revisited in Hunt-Smith et al. (2024), and a similar observation of the positivity of $x\Delta g(x)$ was found. In the present work, using the physics-informed generative ML algorithms, we have solidified our previous results in Khan et al. (2023) with higher precision and up to a larger value of $\omega\sim 14$. While caution is needed regarding the sign of $x\Delta g(x)$ in the small-$x$ region, where LQCD data is not sensitive to $x\Delta g(x)$, the PDFs determined in Fig. 18 show a positive distribution in the interval $x\in[0.05,0.95]$. A recent phenomenological calculation Chakrabarti et al. (2023) has also determined $x\Delta g(x)$ to be positive. In fact, this finding is also consistent with the LQCD calculation Yang et al. (2017) at the physical pion mass, continuum, and infinite volume limits. The LQCD calculation in Yang et al. (2017) obtained the $x$-integrated value of $\Delta G=0.251(47)(16)$ using a local matrix element Ji et al. (2013).
@@
As noted in the previous section, the $xg(x)$ falls off much more slowly than the NNPDF result at large $x$. This may be due to the LQCD calculation being performed at a heavier pion mass and a coarse lattice spacing. Additionally, we observe that because of the large uncertainty in the LQCD matrix elements, the ML algorithms could not isolate any $z$-dependence from the LQCD matrix elements and used data up to $z=0.56$ fm (unlike the polarized data, where Eq. (7) required the use of the LQCD data up to $z=0.36$ fm). As discussed in Sec. V, the large $z$ data used in the ML-generated ITD might be affected by the higher-twist contaminations, leading to a slower fall-off of the PDF. However, as mentioned earlier, the goal of this work is to demonstrate how ML can open an avenue to constrain $xg(x)$ distribution in the mid-to-large $x$ region using LQCD calculations. Future precise LQCD matrix elements will enable isolating the $z$-dependence as an input feature in the ML analyses to obtain $xg(x)$ in the physical and continuum limits. In addition, physics-informed ML application on the LQCD data can have a significant impact on extracting the gluon generalized parton distributions (GPDs). These prospects have been discussed in Kriesten et al. (2022); Allaire et al. (2024); Liuti et al. (2024).
@@ VIII Conlusion and outlook
This work also demonstrates that LQCD nonlocal spatial matrix elements for the gluon helciity distribution beyond $0.36$ fm may contain significant higher-twist contamination, which can impact the PDF determination and should be avoided to ensure the applicability of factorizing the LQCD matrix elements into perturbative Wilson coefficients and nonperturbative PDFs with controllable power corrections. To address this issue, one can train the ML algorithms to learn from short-distance correlations at $z\lesssim 0.36$ fm, thereby minimizing possible contamination from higher-twist effects $\mathcal{O}(\Lambda_{\rm QCD}^{2}z^{2})$ at large spatial distances for a successful reconstruction of the PDFs. With future precise lattice QCD matrix elements, we expect, ML can estimate the correlations between all different $z$ and $p_{z}$ data points and paramterize the leading power corrections, such as $\mathcal{O}(z^{2}\Lambda^{2})$ in the short distance factorization and $\mathcal{O}(\Lambda^{2}/(xp_{z})^{2}),~{}\mathcal{O}(\Lambda^{2}/((1-x)p_{z})^{2})$ in the LaMET. Looking forward, the demonstrated interface between LQCD and ML holds promising potential for investigating the higher-twist contributions at the level of LQCD correlation functions when precise data are available using $z$, $p_{z}$, and $\omega$ as input features in the ML, an area that remains largely unexplored.